OpenStudy (anonymous):

Does it take 3 minutes? While testing a new vaccine, a lab technician noticed that the number of bacteria reduced by half every minute. How long does it take for 3,000 bacteria to reduce to 375?

OpenStudy (anonymous):

oooh every HALF minute sorry

OpenStudy (anonymous):

\[3000\times \left(\frac{1}{2}\right)^{2x}=375\]

OpenStudy (anonymous):

takes one and a half minutes, not 3

OpenStudy (anonymous):

it's reduced by half every MINUTE @satellite73